A fuel is analyzed and found to contain; 75.0 wt% methane, 10.0% ethane, 5.0% ethylene and the balance water. The flowrate of the fuel is 102.7 kg/h. Calculate the molar composition of this gas on both a wet and a dry basis and the ratio(mol H2O/ mol dry gas). If tgis fuel is to be burned with 28% excess air, what is the required air feed rate (kmol/h)? How would the answer change if the combustion were only 40% complete?



1
Expert's answer
2021-04-14T07:46:56-0400

Assuming 100 g of gas, calculate the moles of each component and their mole fractions with a wet basis (with water) and a dry basis (without water) :

n(CH4) = 4.67 mol

n(C2H6) = 0.332 mol

n(C2H4) = 0.178 mol

n(H2O) = 0.544 mol

n(wet) = 5.724 mol

n(dry) = 5.724 - 0.544 = 5.18

Ratio = n(H2O)/n(dry) = 0.544/5.18 = 0.105

To find the ratio of water to dry moles of gas we write the balanced equations for the combustion occuring:

CH4 + 2O2 = CO2 + 2H2O

C2H6 + 3.5O2 = 2CO2 + 3H2O

C2H4 + 3O2 = 2CO2 + 2H2O